OpenStudy (sleepyjess):

@phi can you help on one more? Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (phi):

I would square both equations

OpenStudy (phi):

and try to use cos^2 + sin^2 = 1 to get rid of the sin and cos

OpenStudy (sleepyjess):

Oh gosh, this goes back to trig identities doesn't it?

OpenStudy (phi):

yes, when you see trig functions, expect to use trig identities. but if you use polar coords, you get use to the idea that r^2 = x^2 + y^2 and x = r cos theta , y = r sin theta If that does not make sense, ignore it.

OpenStudy (sleepyjess):

\(x = 3 cos t\\y = 3 sin t\) \(x^2 = 9 cos^2 t\\y^2 = 9 sin^2 t\) \(x^2/9 = cos^2t\\y^2/9 = sin^2t\) \((x/3)^2 + (y/3)^2 = 1\)

OpenStudy (phi):

yes. because that is the equation of a circle, the standard way is to write it x^2 + y^2 = 9

OpenStudy (sleepyjess):

So is that it?

OpenStudy (phi):

your way is fine, but you might see it as I posted it.

OpenStudy (sleepyjess):

Ok :) Thank you!